# Mean field equation and Green’s function on the Tate curve

### 16 Mar 2026 v1

@@ 3.1 Structure of the solution
When $\rho\leq\begin{cases}\frac{p-2}{p}(1-p^{-m}),\ &p>2\\ \frac{1-2^{-m}}{4},&p=2\end{cases}$, then the solution is radial symmetric and monotone.
@@ 3.2 Uniqueness of the solution
In this part, we discuss the uniqueness of the solution. Numerical experiments indicates that for large $\rho$, solutions may not unique, and we can get some solutions that break the properties of radial symmetric. Then by Lemma 3.4 and Lemma 3.5, we can construct many non-radial symmetric solutions. A natural question is whether uniqueness holds under the conditions of $\rho$ guaranteeing radial symmetric of the solution. The answer is affirmative with a slighter stricter condition on $\rho$.
Uniqueness is straightforward on the finite quotient. If $\boldsymbol{u}$ and $\boldsymbol{v}$ both solve (3.4). Then
$\boldsymbol{A}(\boldsymbol{u}-\boldsymbol{v})+\rho(e^{\boldsymbol{u}}-e^{\boldsymbol{v}})=\boldsymbol{0}$
$(\boldsymbol{A}+\rho\text{diag}e^{\boldsymbol{\xi}})(\boldsymbol{u}-\boldsymbol{v})=\boldsymbol{0}$
Lemma 3.9 gives $e^{\xi_{i}}\leq\frac{p}{p-2}$, which is independent of $\rho$. Since $\boldsymbol{A}$ is semi-negative definite, if $\rho\frac{p}{p-2}$ does not exceed the first non-zero eigenvalue of $\boldsymbol{A}$, then $(\boldsymbol{A}+\rho\text{diag}e^{\boldsymbol{\xi}})$ is invertible, then we get the uniqueness.
So the next step is to study the eigenvalue of $\boldsymbol{A}$, i.e. the spectrum of $D$ on the finite quotient $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$. As we can pull back the functions from $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ to $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$, so the eigenvalue of $\boldsymbol{A}$ is just the subset of the spectrum of $D$ by taking the first $d$-th eigenvalues from each component. Alternatively, we can compute the eigenvalues of $\boldsymbol{A}$ via linear algebra as follows:
@@
As $\boldsymbol{A}=p^{-d}\boldsymbol{B}$, we first consider $\boldsymbol{B}$. Denoting $n=(p-1)p^{d-1}$. Then $\boldsymbol{B}\in\mathbb{R}^{mn\times mn}$. From previous observation, $\boldsymbol{B}$ is a block matrix and each block is a circulant matrix. Furthermore, the off-diagonal part is a matrix with the same element. Denoting $\boldsymbol{P}\in\mathbb{R}^{m\times m}$ with $(\boldsymbol{P})_{ij}=p^{-|i-j|}$. And $\boldsymbol{T}\in\mathbb{R}^{n\times n}$ be the circulant part on diagonal with $(\boldsymbol{T})_{ij}=(\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}-1)(1-\delta_{zx})$ (i.e. $\boldsymbol{T}$ is zero on diagonal and $\boldsymbol{T}$ is independent of $v_{p}(z)=v_{p}(x)$). Then we can write $\boldsymbol{B}$ as
$\boldsymbol{B}=\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T}-\text{diag}((\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{1}_{mn})$
@@
$=$ $n\boldsymbol{P}\boldsymbol{1}_{m}\otimes\boldsymbol{1}_{n}+\boldsymbol{1}_{m}\otimes\boldsymbol{T}\boldsymbol{1}_{n}$
@@
$=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}+t_{0}\boldsymbol{I}_{mn}$
@@
$\boldsymbol{B}=\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T}-(n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}+t_{0}\boldsymbol{I}_{mn})$
@@
$=$ $\begin{pmatrix}n\boldsymbol{P}&\\ &\boldsymbol{O}_{mn-m}\end{pmatrix}$
$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}\boldsymbol{T}\boldsymbol{F}=\boldsymbol{I}_{m}\otimes\boldsymbol{\Lambda}$
@@
$=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}$
$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(t_{0}\boldsymbol{I}_{mn})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=t_{0}\boldsymbol{I}_{mn}$
@@
$=$ $\begin{pmatrix}n\boldsymbol{P}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})&\\ &(t_{i}-t_{0})\boldsymbol{I}_{m}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\end{pmatrix}$
@@
$=$ $\frac{1+p-p^{-j}-p^{j+1-m}}{p-1}$
So the eigenvalue of $\boldsymbol{B}$ is $n\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))=(p-1)p^{d-1}\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$ and $t_{i}-t_{0}-ns_{j}=t_{i}-t_{0}-p^{d-1}(1+p-p^{-j}-p^{j+1-m})$, where $i=1,\dots,n-1;\ j=0,\dots,m-1$. Hence, the eigenvalue of $\boldsymbol{B}$ is $(1-p^{-1})\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$ and $p^{-d}(t_{i}-t_{0})-\frac{1+p-p^{-j}-p^{j+1-m}}{p}$. Note that $\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})$ is symmetric, then there exists an orthogonal matrix $\boldsymbol{Q}\in\mathbb{R}^{m\times m}$ to diagonalize it. So the corresponding eigenvector is $\boldsymbol{I}_{m}\otimes\boldsymbol{F}\begin{pmatrix}\boldsymbol{Q}&\\ &\boldsymbol{I}_{m(p-1)p^{d-1}-m}\end{pmatrix}$. Moreover, the spectrum of $\boldsymbol{A}$ must be compatible with the spectrum of $D$. So by Theorem 2.4 we have $p^{-d}(t_{i}-t_{0})-\frac{1+p-p^{-j}-p^{j+1-m}}{p}=-p^{i-1}-p^{i-2}+\frac{p^{-j}+p^{j+1-m}}{p}$. In summary, we get
@@
The smallest non-zero eigenvalue of $\boldsymbol{A}$ is $\begin{cases}(1-p^{-1}),&m=1\\ -(1-p^{-1})\lambda_{1},&m>1\end{cases}$
@@
$\mu_{1}=\text{min}_{\begin{subarray}{c}\boldsymbol{x}\neq\boldsymbol{0}_{m}\\ \boldsymbol{x}\bot\boldsymbol{1}_{m}\end{subarray}}\frac{\boldsymbol{x}^{\top}\boldsymbol{L}\boldsymbol{x}}{\boldsymbol{x}^{\top}\boldsymbol{x}}$
@@
with equality if and only if $p=2$.
@@
$\text{min}_{\begin{subarray}{c}1\leq k\leq m-1\\ 1\leq i\leq d-1\\ 0\leq j\leq m-1\end{subarray}}\{-(1-p^{-1})\lambda_{k},\ p^{i-1}+p^{i-2}-\frac{p^{-j}+p^{j+1-m}}{p}\}$
$=$ $\text{min}\{-(1-p^{-1})\lambda_{1},1-p^{-m}\}$
$=$ $\begin{cases}(1-p^{-1}),&m=1\\ -(1-p^{-1})\lambda_{1},&m>1\end{cases}$
@@
When $\rho\leq\begin{cases}(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),\ &p>2\\ \frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),&p=2\end{cases}$, then the solution of (3.2) is unique.
@@
By Lemma 3.9, when $p>2$, we have
$\rho\frac{p}{p-2}\leq(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$
thus $\rho\leq(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$.
When $p=2$,we have
$4\rho\leq\frac{1}{2}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$
thus $\rho\leq\frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$. ∎
@@ 4 Mean field equation on the Tate curve
Having studied the solution on finite quotient, we now consider the solution on the Tate curve. Since the finite quotient converges to the Tate curve when $d\rightarrow\infty$, we investigate whether the solution converge accordingly. The answer is affirmative. Moreover, the limiting solution retains properties like uniqueness and monotonicity.

### 08 Apr 2026 v2

@@ 3.1 Structure of the solution
When $\rho\in(0,\rho^{*})$, the radially symmetric solution is monotone with a universal upper bound $M$ independent of $d$ and $\rho$, where $\rho^{*}=\theta_{p,m}e^{-M}$ and $\theta_{p,m}$ is the smallest nonzero eigenvalue of $\boldsymbol{A}$. Furthermore, when $\rho\in(0,(1-p^{-m})e^{-M}]$, the solution $u$ must be radially symmetric.
Proof.
Denoting
[…]
and
[…]
First, from Theorem 3.8, we have $(0,p^{-d}-p^{-d-m}]\in S$.
Next, suppose $\rho_{0}\in S$, and the corresponding radially symmetric solution is $\boldsymbol{u}_{0}$. Then by the same argument of Lemma 3.12, we have
$u_{0}(i)\leq M<\begin{cases}\ln\left(\frac{p}{p-2}\right),\ &p>2;\\ \ln 4,&p=2,\end{cases}$
which means any radially symmetric solution have a universal upper bound, so $M$ is well defined. Now considering
$F:\mathbb{R}\times V_{rad}$ $\to V_{rad}$
$(\rho,\boldsymbol{u})$ $\mapsto F(\rho,\boldsymbol{u})=\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}-\rho p^{d}\boldsymbol{e}_{y},$
[…]
$L_{\rho_{0}}:=D_{\boldsymbol{u}}F(\rho_{0},\boldsymbol{u}_{0})=\boldsymbol{A}+\rho_{0}\text{diag}(e^{\boldsymbol{u}_{0}}):V_{rad}\to V_{rad}$
is nondegenerate since $\rho_{0}e^{u_{0}(i)}<\rho^{*}e^{M}=\theta_{p,m}$, then by Implicit Function Theorem, there exists $\varepsilon>0$ s.t.
$(\rho_{0}-\varepsilon,\rho_{0}+\varepsilon)$ $\to V_{rad}$
$\rho$ $\mapsto\boldsymbol{u}_{\rho}$
satisfies $F(\rho,\boldsymbol{u}_{\rho})=0$ and $\boldsymbol{u}_{\rho_{0}}=\boldsymbol{u}_{0}$. Since we can take $\varepsilon$ small enough to ensure $\boldsymbol{u}_{\rho}$ is radially symmetric, so $S$ is open in $(0,\rho^{*})$.
[…]
$f_{s}(u_{l}(c_{0},\dots,c_{s+1}))>f_{s}(u_{l}(c_{0},\dots,c_{s})),$
a contradiction. So $S$ is closed in $(0,\rho^{*})$.
In conclusion, $S=(0,\rho^{*})$. Then by the same argument of Theorem 3.7, we obtain when $\rho\in(0,(1-p^{-m})e^{-M}]$, the solution $u$ must be radially symmetric. ∎
@@ 3.2 Uniqueness of the solution
In this part, we discuss the uniqueness of the solution. Numerical experiments indicates that for large $\rho$, solutions may not unique, and we can get some solutions that break the properties of radially symmetric. Then by Lemma 3.4 and Lemma 3.6, we can construct many non-radially symmetric solutions. Besides, the monotonicity of the solutions may not be ensured for large $\rho$. For instance, if we take $p=3$, $m=1$, $d=2$, $\rho=5$ and $y=1$, then there exists non-monotone solution $\boldsymbol{u}_{nmon}$ and non-radially symmetric solution $\boldsymbol{u}_{nrad}$, see in Table 1.
$\boldsymbol{u}_{nmon}$ $\boldsymbol{u}_{nrad}$
$u(1)$ -18.2689 -22.2707
$u(2)$ -1.3831 -0.4886
$u(4)$ 1.4168 1.9684
$u(5)$ -1.3831 -0.4886
$u(7)$ 1.4168 -8.7704
$u(8)$ -1.3831 -0.4886
Table 1: Non-monotone solutions and non-radially symmetric solutions
A natural question is whether uniqueness holds under the conditions of $\rho$ guaranteeing radially symmetric and monotonicity of the solution, the answer is affirmative. In order to investigate it, our first step is to study the eigenvalue of $\boldsymbol{A}$, i.e. the spectrum of $D$ on the finite quotient $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$. As we can pull back the functions from $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ to $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$, so the eigenvalue of $\boldsymbol{A}$ is just the subset of the spectrum of $D$ by taking the first $d$-th eigenvalues from each component. Alternatively, we can compute the eigenvalues of $\boldsymbol{A}$ via linear algebra as follows:
@@
As $\boldsymbol{A}=p^{-d}\boldsymbol{B}$, we first consider $\boldsymbol{B}$. Denoting $n=(p-1)p^{d-1}$. Then $\boldsymbol{B}\in\mathbb{R}^{mn\times mn}$. From previous observation, $\boldsymbol{B}$ is a block matrix and each block is a circulant matrix. Furthermore, the off-diagonal part is a matrix with the same element. Denoting $\boldsymbol{P}\in\mathbb{R}^{m\times m}$ with $(\boldsymbol{P})_{ij}=p^{-|i-j|}$. And $\boldsymbol{T}\in\mathbb{R}^{n\times n}$ be the circulant part on diagonal with $(\boldsymbol{T})_{ij}=\left(\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}-1\right)(1-\delta_{zx})$ (i.e. $\boldsymbol{T}$ is zero on diagonal and $\boldsymbol{T}$ is independent of $v_{p}(z)=v_{p}(x)$). Then we can write $\boldsymbol{B}$ as
$\boldsymbol{B}=\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T}-\text{diag}((\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{1}_{mn}).$
@@
$=$ $n\boldsymbol{P}\boldsymbol{1}_{m}\otimes\boldsymbol{1}_{n}+\boldsymbol{1}_{m}\otimes\boldsymbol{T}\boldsymbol{1}_{n},$
@@
$=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}+t_{0}\boldsymbol{I}_{mn},$
@@
$\boldsymbol{B}=\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T}-(n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}+t_{0}\boldsymbol{I}_{mn}).$
@@
$=$ $\begin{pmatrix}n\boldsymbol{P}&\\ &\boldsymbol{O}_{mn-m}\end{pmatrix}$,
$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}\boldsymbol{T}\boldsymbol{F}=\boldsymbol{I}_{m}\otimes\boldsymbol{\Lambda},$
@@
$=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n},$
$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(t_{0}\boldsymbol{I}_{mn})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=t_{0}\boldsymbol{I}_{mn}.$
@@
$=$ $\begin{pmatrix}n\boldsymbol{P}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})&\\ &(t_{i}-t_{0})\boldsymbol{I}_{m}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\end{pmatrix}$.
@@
$=$ $\frac{1+p-p^{-j}-p^{j+1-m}}{p-1}.$
So the eigenvalue of $\boldsymbol{B}$ is
$n\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))=(p-1)p^{d-1}\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$
and
$t_{i}-t_{0}-ns_{j}=t_{i}-t_{0}-p^{d-1}(1+p-p^{-j}-p^{j+1-m}),$
where $i=1,\dots,n-1;\ j=0,\dots,m-1$. Hence, the eigenvalue of $\boldsymbol{A}$ is
$(1-p^{-1})\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$
and
$p^{-d}(t_{i}-t_{0})-\frac{1+p-p^{-j}-p^{j+1-m}}{p}.$
Note that $\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})$ is symmetric, then there exists an orthogonal matrix $\boldsymbol{Q}\in\mathbb{R}^{m\times m}$ to diagonalize it. So the corresponding eigenvector is $\boldsymbol{I}_{m}\otimes\boldsymbol{F}\begin{pmatrix}\boldsymbol{Q}&\\ &\boldsymbol{I}_{m(p-1)p^{d-1}-m}\end{pmatrix}$. Moreover, the spectrum of $\boldsymbol{A}$ must be compatible with the spectrum of $D$. So by Theorem 2.4 we have $p^{-d}(t_{i}-t_{0})-\frac{1+p-p^{-j}-p^{j+1-m}}{p}=-p^{i-1}-p^{i-2}+\frac{p^{-j}+p^{j+1-m}}{p}$. In summary, we get
@@
The smallest non-zero eigenvalue of $\boldsymbol{A}$ is
$\theta_{p,m}=\begin{cases}(1-p^{-1}),&m=1;\\ -(1-p^{-1})\lambda_{1},&m>1.\end{cases}$
@@
$\mu_{1}=\min_{\begin{subarray}{c}\boldsymbol{x}\neq\boldsymbol{0}_{m}\\ \boldsymbol{x}\bot\boldsymbol{1}_{m}\end{subarray}}\frac{\boldsymbol{x}^{\top}\boldsymbol{L}\boldsymbol{x}}{\boldsymbol{x}^{\top}\boldsymbol{x}}.$
@@
with equality if and only if $p=2$, $\forall m$ and $m=1$, $\forall p$.
@@
$\theta_{p,m}=$ $\min_{\begin{subarray}{c}1\leq k\leq m-1\\ 1\leq i\leq d-1\\ 0\leq j\leq m-1\end{subarray}}\left\{-(1-p^{-1})\lambda_{k},\ p^{i-1}+p^{i-2}-\frac{p^{-j}+p^{j+1-m}}{p}\right\}$
$=$ $\min\{-(1-p^{-1})\lambda_{1},1-p^{-m}\}$
$=$ $\begin{cases}(1-p^{-1}),&m=1;\\ -(1-p^{-1})\lambda_{1},&m>1.\end{cases}$
@@
When $\rho\in(0,\rho^{*})$, then the solution of (3.2) is unique.
@@
As $\theta_{p,m}\leq(1-p^{-m})$ by Lemma 3.15, then the solution must be radially symmetric. If $\boldsymbol{u}$ and $\boldsymbol{v}$ both solve (3.4). Then
$\boldsymbol{A}(\boldsymbol{u}-\boldsymbol{v})+\rho(e^{\boldsymbol{u}}-e^{\boldsymbol{v}})=\boldsymbol{0}.$
$(\boldsymbol{A}+\rho\text{diag}e^{\boldsymbol{\xi}})(\boldsymbol{u}-\boldsymbol{v})=\boldsymbol{0}.$
Theorem 3.13 gives $e^{\xi_{i}}\leq e^{M}$, which is independent of $\rho$ and $d$. Besides, as $\boldsymbol{A}$ is semi-negative definite and
$\rho e^{\xi_{i}}<\rho^{*}e^{M}=\theta_{p,m},$
then $\rho e^{\xi_{i}}$ does not exceed the first non-zero eigenvalue of $\boldsymbol{A}$. Then $(\boldsymbol{A}+\rho\text{diag}e^{\boldsymbol{\xi}})$ is invertible, then we get the uniqueness. ∎
Remark 3.17.
The order of $\rho^{*}$ is optimal. Since $\sum_{i}e^{u_{i}}=p^{d}$, we have $m(p-1)p^{d-1}e^{M}\geq p^{d}$, therefore $M>\ln\left(\frac{p}{m(p-1)}\right)$, which means
$\rho^{*}<\frac{m(p-1)}{p}(1-p^{-m}).$
On the other hand, since
$M<\begin{cases}\ln\left(\frac{p}{p-2}\right),\ &p>2;\\ \ln 4,&p=2,\end{cases}$
which means
$\rho^{*}>\begin{cases}\frac{1}{4}\theta_{2,m},&p=2;\\ (1-2p^{-1})\theta_{p,m},\ &p>2.\end{cases}=\begin{cases}\frac{1}{8},&p=2,\ m=1;\\ -\frac{1}{8}\lambda_{1},&p=2,\ m>1;\\ (1-2p^{-1})(1-p^{-1}),\ &p>2,\ m=1;\\ -(1-2p^{-1})(1-p^{-1})\lambda_{1},&p>2,\ m>1.\end{cases}$
@@ 4 Mean field equation on the Tate curve
Since the finite quotient converges to the Tate curve when $d\rightarrow\infty$, we investigate whether the solution converge accordingly. From the numerical experiments, we observe that the solution on the finite quotient is convergent. Moreover, the limiting solution retains properties like uniqueness and monotonicity.